If 3.6 L of HF decomposed at 65°C and 1.35 atm, what mass of SnF2 will be produced given the reaction:

Sn (s)+2 HF (g)———SnF2 (s)+H2 (g)


1
Expert's answer
2022-04-15T03:04:02-0400

Sn + 2HF = SnF2 + H2


PV=nRT

P =  1.35 atm, V = 3.6 L, T = 65°C = 338.15 K, R = 0.0821 atm * L / mol * K


n(HF) = PV/RT, n(HF) = (1.35 atm * 3.6 L) / (0.0821 atm*L /mol*K * 338.15 K) = 17.51 moles


by equation

m(SnF2) = (m(HF) * Mr(SnF2)) / 2Mr(HF) = (n(HF)/2) * Mr(SnF2) = 17.51/2 * 157 = 1374.54 g